hartnn (hartnn):

\[\begin{array}{|c|c|c|c|c|}\hline Sr no.&A&B&D&E\\ \hline T&T&T&T&T\\T&F&F&T&F\\F&T&T&F&F\\F&F&T&T&T\\ \hline\end{array}\]

hartnn (hartnn):

\[\begin{array}{|c|c|c|c|c|}\hline Sr \:no.&A&B&D&E\\ \hline 1&T&T&T&T\\T&F&F&T&F\\F&T&T&F&F\\F&F&T&T&T\\ \hline\end{array}\] just trying to insert table in latex

hartnn (hartnn):

\[\begin{array} &Y&X \\ &W&Z\end{array}\] `\[\begin{array} &Y&X \\ &W&Z\end{array}\]` not aligned

hartnn (hartnn):

\[\begin{array} {|c|c|c|c|c|} &Y&X \\ &W&Z\end{array}\] `\[\begin{array} {|c|c|c|c|c|} &Y&X \\ &W&Z\end{array}\]`

hartnn (hartnn):

4 lines with {|c|c|c|c|c|} how ?

hartnn (hartnn):

\[\begin{array} &&Y&X \\ &W&Z\end{array}\] `\[\begin{array} &&Y&X \\ &W&Z\end{array}\]`

hartnn (hartnn):

\[\begin{array}{|c|c|}\hline Expression&Substitution\\ \hline \sqrt{a^2-x^2}&x=asin \theta \quad or \quad x=acos\theta \\ \hline \sqrt{x^2-a^2}&F\\ \hline F&T\\ \hline F&F\\ \hline\end{array}\] `\[\begin{array}{|c|c|}\hline Expression&Substitution\\ \hline \sqrt{a^2-x^2}&x=asin \theta \quad or \quad x=acos\theta \\ \hline \sqrt{x^2-a^2}&F\\ \hline F&T\\ \hline F&F\\ \hline\end{array}\]`

hartnn (hartnn):

\[\begin{array}{|cc|}1&1&1&1&1&1&1&1&1\end{array}\] `\[\begin{array}{|cc|}1&1&1&1&1&1&1&1&1\end{array}\]` hmm.. \[\begin{array}{|c|c|}1&1&1&1&1&1&1&1&1\end{array}\] `\[\begin{array}{|c|c|}1&1&1&1&1&1&1&1&1\end{array}\]` \[\begin{array}{|c|c}1&1&1&1&1&1&1&1&1\end{array}\] `\[\begin{array}{|cc|}1&1&1&1&1&1&1&1&1\end{array}\]` \[\begin{array}{c|c}1&1&1&1&1&1&1&1&1\end{array}\] `\[\begin{array}{c|c}1&1&1&1&1&1&1&1&1\end{array}\]`

hartnn (hartnn):

\[\hline \] that can only be used inside \begin {array} and end {array} ? \[|c|c\] that can only be used inside \begin {array} and end {array} ?

hartnn (hartnn):

\[\begin{array} Y&X \\ W&Z\end{array}\] `\[\begin{array} Y&X \\ W&Z\end{array}\]` \[\begin{array} &Y&X \\ W&Z\end{array}\] `\[\begin{array} &Y&X \\ W&Z\end{array}\]` \[\begin{array} %Y&X \\ W&Z\end{array}\]

hartnn (hartnn):

`\begin{array} %Y&X \\ W&Z\end{array}`

hartnn (hartnn):

so any character is required after \begin{array} and which is not displayed.....

hartnn (hartnn):

\[\begin{array} AY&X \\ W&Z\end{array}\] `\[\begin{array} AY&X \\ W&Z\end{array}\]`

hartnn (hartnn):

how do i insert line between 2 statements ? \(\text Line 1. \\ \text{*I want to insert a horizontal line here.*}\\Line2.\)

hartnn (hartnn):

\[\begin{array} & \\ \hline & \end{array} \begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\] `\[\begin{array} & \\ \hline & \end{array} \begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\begin{array} & \\ \hline & \end{array}\]` I tried this :P there must be a better way...........

hartnn (hartnn):

\( \\ \_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_ \) continuous line ??

hartnn (hartnn):

^ `\\ \_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_`

OpenStudy (accessdenied):

Hmm... so you want to separate two different sections outside of an array? Like a section on trig integrals, divide with a line, and then start logarithmic integrals?

hartnn (hartnn):

that is correct.

hartnn (hartnn):

i would show u what i did... \[ \huge \text{In general,}\int \frac{f’(x)}{f(x)}\:dx=\ln|f(x)|+c \\ \_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_ \\ \huge \quad \quad \quad \quad \color{blue}{\text{Trigonometric Integrals}} \\ \huge 5. \int \sin\: x\:dx=-\cos\:x+c \]

hartnn (hartnn):

O.o does that look continuous line now ? when i posted it , first it looked like _ _ _ `\huge \text{In general,}\int \frac{f’(x)}{f(x)}\:dx=\ln|f(x)|+c \\ \_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_ \\ \huge \quad \quad \quad \quad \color{blue}{\text{Trigonometric Integrals}} \\ \huge 5. \int \sin\: x\:dx=-\cos\:x+c `

OpenStudy (accessdenied):

I'm not sure if there's a good way to do that in the mathjax setup here. There is a command for it that I see in a full LaTeX document, but it is not applicable in the math environment that you work in on here... Instead, I would probably just space out separate sections a bit and use a larger title that stands out from paragraph text to provide.more of a contrasting feel. -- I think you'll have different effects between inline-set LaTeX and block-set LaTeX. Like, using \ ( and \ ) v.s. using \ [ and \ ]

hartnn (hartnn):

then can u tell how to insert a blank line or good amount of space between 2 lines ? \( \huge \text{In general,}\int \frac{f’(x)}{f(x)}\:dx=\ln|f(x)|+c \\ \text{i want some space here} \\ \huge \quad \quad \quad \quad \color{blue}{\text{Trigonometric Integrals}} \\ \huge 5. \int \sin\: x\:dx=-\cos\:x+c \)

OpenStudy (accessdenied):

\[ a \\ ~ \\ ~ \\ ~\\ b \] I can't remember the code tag here... :P Math environment doesn't like empty lines, so I just add a tilde and then a second line break. The tilde acts as a space and then it interprets it as its own line. Of course, you may also break the LaTeX into multiple sections if you like.

hartnn (hartnn):

like how would i 'just space out separate sections a bit'

hartnn (hartnn):

okk..

hartnn (hartnn):

\(\huge \text{In general,}\int \frac{f’(x)}{f(x)}\:dx=\ln|f(x)|+c \\~ \\~ \\~ \\~ \\ \huge \quad \quad \quad \quad \color{blue}{\text{Trigonometric Integrals}} \\ \huge 5. \int \sin\: x\:dx=-\cos\:x+c\)

hartnn (hartnn):

thanks!

OpenStudy (accessdenied):

You're welcome. :)

hartnn (hartnn):

one last thing , 'There is a command for it that I see in a full LaTeX document,' <---which command ?

OpenStudy (accessdenied):

I saw some things like \hrulefill and \rule in some posts when I was google'ing. When I tried them out though, they didn't work. :P

hartnn (hartnn):

okay :)